Calculate 250% of 100 and calculate 100% of 250.
What do you notice about your answers. Explain your answers.
1
Expert's answer
2019-05-06T04:29:58-0400

250% of 100:

"100\cdot\frac{250\%}{100\%}=250"

100% of 250:


"250\cdot\frac{100\%}{100\%}=250"

The both 250% of 100 and 100% of 250 give the same value. As we can notice, x% of 100 gives x%. On another hand, 100% of any value gives this value.